# Chain-of-Thought Unfaithfulness as Disguised Accuracy

### v1

@@ Abstract
Understanding the extent to which Chain-of-Thought (CoT) generations align with a large language model’s (LLM) internal computations is critical for deciding whether to trust an LLM’s output. As a proxy for CoT faithfulness, Lanham et al. (2023) propose a metric that measures a model’s dependence on its CoT for producing an answer. Within a single family of proprietary models, they find that LLMs exhibit a scaling-then-inverse-scaling relationship between model size and their measure of faithfulness, and that a 13 billion parameter model exhibits increased faithfulness compared to models ranging from 810 million to 175 billion parameters in size. We evaluate whether these results generalize as a property of all LLMs. We replicate their experimental setup with three different families of models and, under specific conditions, successfully reproduce the scaling trends for CoT faithfulness they report. However, we discover that simply changing the order of answer choices in the prompt can reduce the metric by 73 percentage points. The faithfulness metric is also highly correlated ($R^{2}$ = 0.91) with accuracy, raising doubts about its validity as a construct for evaluating faithfulness.
@@ 1 Introduction
  • RQ2: If so, do all LLMs start to produce unfaithful CoTs at a similar model size (13B parameters)?
@@
We measure the faithfulness of three families of openly accessible LLMs — Llama 2 (Touvron et al., 2023), Flan-T5 (Chung et al., 2022) + Flan-UL2 (Tay et al., 2023), and Pythia (Biderman et al., 2023) DPO — using Lanham et al. (2023)’s metric on the same set of multiple choice benchmarks and addition tasks.
Initially, our results show similar scaling trends, with unfaithfulness increasing for models above a certain capability level. However, after randomly shuffling the answer choices so that they appear in different orderings for the with- and without-CoT settings, the high unfaithfulness scores for less capable models become low, dropping by as much as 73 percentage points in some cases. We discuss the implications of this finding with respect to measuring the faithfulness of CoTs and hypothesize another possible mechanism by which it could be confounded—the ability of larger models to capture the knowledge of the CoTs in their weights. Although we are able to replicate these scaling trends for other LLMs under certain circumstances, we also find strong linear correlation between the faithfulness metric and accuracy, bringing into question how informative the metric is altogether.
@@ 2 Related Work
With the training of increasingly large models (measured by number of parameters) has come an interest in deriving scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022). While many tasks exhibit a trend of increased performance with larger model sizes, other trends such as inverse scaling (Lin et al., 2022; McKenzie et al., 2022) and U-shaped scaling (Belkin et al., 2018; Black et al., 2022; Ruis et al., 2023; Wei et al., 2023) also occur. Lanham et al. (2023)’s results can be viewed as an instance of inverse-scaling with respect to their measure of faithfulness for models larger than 13b. In our study, we ask whether the inverse scaling trend of their metric occurs for LLMs generally.
@@ 3 Experimental Setup
In this section, we describe the metric, models, tasks, and implementation details we followed. We indicate where our methods align with Lanham et al., and we motivate our alternative approaches where appropriate.
Measuring Unfaithfulness
Lanham et al. (2023) introduce a metric of unfaithfulness that measures how often a model $\mathcal{M}$ produces the same prediction with- and without-CoT on a given task’s dataset $\mathcal{D}$. Equation 1 below shows this calculation, where NoCoT denotes the answer provided by the model without CoT prompting, and CoT denotes the answer provided by the model with CoT prompting.
@@
  • FLAN-T5 + UL2 (Chung et al., 2022; Tay et al., 2023) is an encoder-decoder family of LLMs available with 77M, 248M, 783M, 2.85B, 11.3B, and 20B parameters. Unlike Anthropic’s models or the Llama 2 family of models, it hasn’t undergone human preference alignment, however it has been instruction finetuned for over 1800 natural language tasks.
  • Pythia DPO extends Pythia pretrained decoder-only LLMs (Biderman et al., 2023), by finetuning on Anthropic’s human preference dataset (Bai et al., 2022) using direct preference optimization (DPO; Rafailov et al., 2023). These models cover 70M, 160M, 410M, 1B, 1.4B and 2.8B parameters. [^1] Links to Pythia DPO Huggingface models: 70M, 160M, 410M, 1B, 1.4B, 2.8B
Figure 2: Illustration of the same vs. different ordering conditions for MCQs (§3). In the same condition, the model is shown the same choice order in the with- and without-CoT settings. In the different condition, the choice ordering is shuffled so that the with- and without-CoT settings are shown a different order.
FLAN-T5 and Pythia DPO don’t cover both sides of the 13B parameter point-of-interest where we’d expect to find the V-shape. However, they do provide context about the relationship between model scaling and CoT reliance for models smaller than 13B parameters. We don’t include other common benchmark models like Mistral (Jiang et al., 2023) or MPT (Team, 2023b; a) because they don’t have the range of model sizes needed to extrapolate patterns related to model scaling.
Ideally, we would include the Anthropic models used in Lanham et al. (2023), taken from Ganguli et al. (2023). We can’t use these models, even through the Anthropic API, because we are restricted to a single model size, specifically, the size of Anthropic’s Claude model available at the moment. We need access to all model sizes in Ganguli et al. (2023) to be able to fully recreate the measurements necessary for drawing conclusions about model scaling. We also can’t guarantee that Anthropic’s current Claude model is any of the models in Lanham et al. (2023) since Anthropic may have updated the publicly available model since the paper was released. Finally, even with access to all model sizes through an API, the decoding strategy described in the original paper requires manual intervention in the decoding process to append special prompts for extracting answers. As a result, we assume the original findings by Lanham et al. (2023) are valid, and instead look to evaluate other model families under the same conditions, where possible, to validate whether their findings generalize to other LLMs. [^2] We use “model family” to mean a set of LLMs that range in parameter count, but share similar pretraining procedures.
Multiple Choice Task
@@
Figure 3: Lanham et al. (2023)’s unfaithfulness (top) and CoT prompting accuracy (bottom) across different model sizes. Left plots show same choice ordering, right plots different ordering (§3). Each point is an average across all benchmarks, and the error bars show the standard deviation across benchmarks.
Prior work has shown that models are biased towards producing certain outputs, even when input has no content, e.g., “Input: N/A Sentiment:” leads to a higher rate of producing “positive” than “negative” (Zhao et al., 2021). Given this context, as well as the fact that TruthfulQA’s validation dataset always has “A” as the correct answer, we employ two strategies to shuffle the ordering of the MCQ choices. These conditions are illustrated and described below. [^3] We also have additional results for each evaluation dataset’s original order, available in Figures 6 and 7 in the Appendix.
Lanham et al. (2023) don’t report details surrounding choice ordering, so evaluating in both setups allows for a fairer comparison. By accounting for confounding factors like the model’s token preference, we also ensure that the different ordering results are isolated to only the content of the choices.
@@
In order to more carefully measure the effect of task complexity on their faithfulness metric, Lanham et al. (2023) propose an addition task where the difficulty is controlled by varying the number of operands involved in the computation. We implement this task as 2-digit and 3-digit addition problems with 2, 4, 8, or 16 operands and prompt the model to provide it’s answer in “<answer></answer>” XML tags. For each digit operands pair we generate a set of 2000 addition problems by sampling uniformly the operands from the appropriate range of integers (i.e. 10-99 for 2-digit and 100-999 for 3-digit). Using the same models as previously, we run inference with and without CoT for each addition task and calculate the same answer percentage.
@@ 4 Results
In this section, we address whether other LLMs exhibit inverse scaling in model size for CoT faithfulness. If so, we aim to determine the parameter count at which the inflection point occurs and whether this depends on task difficulty. Figure 3 shows how unfaithfulness (upper subfigures) and accuracy with CoTs (lower subfigures) change with different model sizes. We refer the reader to Tables 3-11 in the Appendix for a more detailed breakdown of these results.
Figure 4: Lanham et al. (2023)’s unfaithfulness and task accuracy for 2- and 3-digit addition problems containing 2, 4, 8, or 16 operands using Llama 2. The bottom plots show the accuracy for each condition, where CoT prompting accuracy is represented a dashed line, and without-CoT accuracy is represented with a solid line. The x-axis for all plots is a log-scale with the model size (number of model parameters).
@@
Yes. In Figure 3, the same condition (left subfigures), the Pythia DPO models exhibit a V-shaped pattern of CoT unfaithfulness similar to that identified by Lanham et al. (2023) in Figure 1 which flattens in the different condition (right subfigures in Figure 3). The flattening happens because Pythia DPO models seem to favor specific answers over random selection when they are not yet capable of solving a task. Moreover, if we focus on the accuracy lines that are one standard deviation below the mean in lower subfigures of Figure 3, we observe that only Llama 2 models and FLAN-UL2 ever become better than a random baseline for all benchmarks. [^4] A baseline that selects an answer choice at random ranges 20–25% depending on the benchmark. In both the same and different conditions, we see inverse scaling in the Llama 2 unfaithfulness scores; CoTs become less faithful as the models get bigger. These results support the takeaway of Lanham et al. (2023): “only models of a certain capability level (but no higher) on a task seem to produce faithful CoT.”
@@
The observed connection between unfaithfulness and task accuracy leads us to further explore the correlation between these two factors. In our accuracy–unfaithfulness plots, shown in Figure 5, we discover a high correlation in the different condition. We contend that this strikingly clear correlation between unfaithfulness and accuracy with CoTs suggests a simplification of the intricate concept of measuring reasoning faithfulness that is too reductive.
@@
No. Lanham et al. (2023) find that inverse scaling only begins at 13B parameters. The Llama 2 unfaithfulness scores suggest that inverse scaling begins at a model size smaller than 7B parameters. The Pythia DPO unfaithfulness scores in the same ordering condition exhibit a V-shape similar to the one found by Lanham et al. (2023), however, it indicates inverse scaling begins at 1B parameters. In the different condition, the Pythia DPO model has constant unfaithfulness.
Figure 5: Task accuracy vs. Lanham et al. (2023)’s unfaithfulness metric in the same and different MCQ choice ordering settings. The dashed black line corresponds to a linear regression line fit to the data, along with its respective $R^{2}$ correlation value. In the same condition, the correlation between accuracy and unfaithfulness is minimal ($R^{2}=0.059$), but strong ($R^{2}=0.912$) in the different condition.
@@
No. To answer this, we consider Figure 4, which shows the Llama 2 models’ performance on the addition task. Given that 2-digit addition is an easier task than 3-digit addition, we would expect the inverse scaling trend to begin at smaller models for 2- than for 3-digit addition. We don’t see this pattern in the Llama 2 unfaithfulness scores, where both settings point to the 13B model having the most faithful CoTs. The discrepancy in our findings compared to Lanham et al. (2023) may be an issue about granularity, where our three Llama 2 model sizes don’t provide enough detail to see a difference between the addition conditions. Similar to findings in the original paper, we omit results for Pythia DPO and FLAN-T5 due to unreliability in extracting answers from the model’s response for smaller models.
@@ 5 Discussion and Conclusions
In §4, we demonstrate that high unfaithfulness scores by models performing no better than random do not necessarily stem from genuinely unfaithful CoTs but due to the models’ tendency to prefer specific answers. Thus, a more accurate understanding of CoTs faithfulness can be obtained by computing Lanham et al. (2023)’s unfaithfulness measurement in the different choice ordering setting that we use in this work.
@@
Moreover, with the different choice ordering setting, we find a strong linear correlation between task accuracy and unfaithfulness of generated CoTs. We contest the notion that higher accuracy in one LLM compared to another means its CoTs are automatically less faithful. If these CoTs are also regularly plausible — which is often the case with highly capable models like GPT-4 — the correlation suggests that such models reason differently from people in all cases where they solve the task well. Finally, drops as big as 73 percentage points (Pythia DPO 70M) in the unfaithfulness measurement just from changing the choice order raise concerns about the sensitivity of this measurement. Together, these observations caution against using the rate at which answers change with the provision of CoTs as a measurement for assessing their faithfulness.

### v3

@@ Abstract
Understanding the extent to which Chain-of-Thought (CoT) generations align with a large language model’s (LLM) internal computations is critical for deciding whether to trust an LLM’s output. As a proxy for CoT faithfulness, Lanham et al. (2023) propose a metric that measures a model’s dependence on its CoT for producing an answer. Within a single family of proprietary models, they find that LLMs exhibit a scaling-then-inverse-scaling relationship between model size and their measure of faithfulness, and that a 13 billion parameter model exhibits increased faithfulness compared to models ranging from 810 million to 175 billion parameters in size. We evaluate whether these results generalize as a property of all LLMs. We replicate the experimental setup in their section focused on scaling experiments with three different families of models and, under specific conditions, successfully reproduce the scaling trends for CoT faithfulness they report. However, after normalizing the metric to account for a model’s bias toward certain answer choices, unfaithfulness drops significantly for smaller less-capable models. This normalized faithfulness metric is also strongly correlated ($R^{2}$=0.74) with accuracy, raising doubts about its validity for evaluating faithfulness.
@@ 1 Introduction
  • RQ2: If so, do other non-proprietary families of LLMs start to produce unfaithful CoTs at a similar model size (13B parameters)?
@@
We measure the faithfulness of three families of openly accessible LLMs — Llama 2 (Touvron et al., 2023), FLAN-T5 (Chung et al., 2022) + FLAN-UL2 (Tay et al., 2023), and Pythia DPO (O’Mahony et al., 2024) — using Lanham et al. (2023)’s metric on the same set of multiple choice benchmarks and addition tasks.
Initially, our results show similar scaling trends, with unfaithfulness increasing for models above a certain capability level. However, after normalizing the metric to account for a model’s preference for certain answer choices, the high unfaithfulness scores for less capable models become low. That is, the rate at which these models produce the same answer with the addition of CoT is largely accounted for by a bias to select a certain letter option.
We discuss the implications of this finding with respect to measuring the faithfulness of CoTs and hypothesize another possible mechanism by which it could be confounded—the ability of larger models to capture the knowledge of the CoTs in their weights. Although we are able to replicate these scaling trends for other LLMs under certain circumstances, we also find strong linear correlation between the faithfulness metric and accuracy, bringing into question how informative the metric is altogether.
@@ 2 Related Work
With the training of increasingly large models (measured by number of parameters) has come an interest in deriving scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022). While many tasks exhibit a trend of increased performance with larger model sizes, other trends such as inverse scaling (Lin et al., 2022; McKenzie et al., 2022) and U-shaped scaling (Belkin et al., 2019; Black et al., 2022; Ruis et al., 2023; Wei et al., 2023) also occur. Lanham et al. (2023)’s results can be viewed as an instance of inverse-scaling with respect to their measure of faithfulness for models larger than 13b. In our study, we ask whether the inverse scaling trend of their metric occurs for LLMs generally.
@@ 3 Experimental Setup
In this section, we describe the metrics, models, tasks, and implementation details we followed to reproduce the scaling experiments. [^1] We release our code at: https://github.com/utahnlp/cot_disguised_accuracy We indicate where our methods align with Lanham et al., and we motivate our alternative approaches where appropriate.
Lanham et al. Unfaithfulness
To study the relationship between model size and faithfulness, Lanham et al. (2023) use a metric of unfaithfulness that measures how often a model $\mathcal{M}$ produces the same prediction with- and without-CoT on a given task’s dataset $\mathcal{D}$ [^2] Note that Lanham et al. (2023) introduce two additional measures of faithfulness based on early answering and adding mistakes before studying how faithfulness scales with model sizes. However, for the scaling study, they say that the measurement in (1) is “highly predictive of overall early answering and adding mistakes results. […] We [Lanham et al.] thus use this metric in lieu of running the full set of early answering and adding mistakes experiments for computational reasons.” Equation (1) below shows this calculation, where NoCoT denotes the answer provided by the model without CoT prompting, and CoT denotes the answer provided by the model with CoT prompting.
@@
Normalized Lanham et al. Unfaithfulness
Prior work has shown that models are biased towards producing certain outputs, even when the input has no content, e.g., “Input: N/A Sentiment:” leads to a higher rate of producing “positive” than “negative” (Zhao et al., 2021). Given this context, as well as the fact that TruthfulQA’s validation dataset always has “A” as the correct answer, we introduce a normalized version of the unfaithfulness metric which accounts for a model’s tendency to select certain letter choices regardless of the content.
We first compute the normalization term as:
$\textsc{N}(\mathcal{M},\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\mathbbm{1}_{\left[\textsc{NoCoT}(\mathcal{M},x)=\textsc{NoCoT}(\mathcal{M},\tilde{x})\right]}$ (2)
where $\tilde{x}$ is a version of $x$ where the answer choices have been randomly shuffled. In other words, Equation (2) measures how often the model selects the same letter choice when prompted twice in the No-CoT setting with different answer choice orderings.
Then the normalized metric becomes:
$\textsc{Unfaithfulness}_{\text{Normalized}}(\mathcal{M},\mathcal{D})=\frac{\textsc{Unfaithfulness}_{\text{Lanham}}(\mathcal{M},\mathcal{D})}{\textsc{N}(\mathcal{M},\mathcal{D})}$ (3)
This measures the frequency of answer changes with CoT, compared to changes expected from merely shuffling the answer order. [^3] We also experimented with shuffling the ordering of answer choices between the No-CoT and CoT settings. These results are found in Appendix A.2.
@@
  • FLAN-T5 + UL2 (Chung et al., 2022; Tay et al., 2023) is an encoder-decoder family of LLMs available with 77M, 248M, 783M, 2.85B, 11.3B, and 20B parameters. Unlike Anthropic’s models or the Llama 2 family of models, it has not undergone human preference alignment, however it has been instruction finetuned for over 1800 natural language tasks.
  • Pythia DPO (O’Mahony et al., 2024) extends Pythia pretrained decoder-only LLMs (Biderman et al., 2023), by finetuning on Anthropic’s human preference dataset (Bai et al., 2022) using direct preference optimization (DPO; Rafailov et al., 2023). These models cover 70M, 160M, 410M, 1B, 1.4B and 2.8B parameters. [^4] Links to Pythia DPO Huggingface models: 70M, 160M, 410M, 1B, 1.4B, 2.8B.
FLAN-T5 and Pythia DPO do not cover both sides of the 13B parameter point-of-interest where we would expect to find the V-shape. However, they do provide context about the relationship between model scaling and CoT reliance for models smaller than 13B parameters. We do not include other common benchmark models like Mistral (Jiang et al., 2023) or MPT (Team, 2023b; a) because they do not have the range of model sizes needed to extrapolate patterns related to model scaling.
Ideally, we would include the Anthropic models used in Lanham et al. (2023), taken from Ganguli et al. (2023). We cannot use these models, even through the Anthropic API, because we are restricted to a single model size, specifically, the size of Anthropic’s Claude model available at the moment. We need access to all model sizes in Ganguli et al. (2023) to be able to fully recreate the measurements necessary for drawing conclusions about model scaling. We also cannot guarantee that Anthropic’s current Claude model is any of the models in Lanham et al. (2023) since Anthropic may have updated the publicly available model since the paper was released. Finally, even with access to all model sizes through an API, the decoding strategy described in the original paper requires manual intervention in the decoding process to append special prompts for extracting answers. As a result, we assume the original findings by Lanham et al. (2023) are valid, and instead look to evaluate other model families under the same conditions, where possible, to validate whether their findings generalize to other LLMs. [^5] We use “model family” to mean a set of LLMs that range in parameter count, but share similar pretraining procedures.
Multiple Choice Benchmarks
@@
Figure 2: Lanham et al. (2023)’s unfaithfulness, our normalized unfaithfulness, and CoT prompting accuracy across different model sizes for Pythia DPO and Llama 2 model families; see §3 for details. Each point is a model, corresponding to a model family (symbol) and size, evaluated on a given benchmark (color).
@@
In order to more carefully measure the effect of task complexity on their faithfulness metric, Lanham et al. (2023) propose an addition task where the difficulty is controlled by varying the number of operands involved in the computation. We implement this task as 2-digit and 3-digit addition problems with 2, 4, 8, or 16 operands and prompt the model to provide its answer in “<answer></answer>” XML tags. For each digit operands pair we generate a set of 2000 addition problems by sampling uniformly the operands from the appropriate range of integers (i.e. 10-99 for 2-digit and 100-999 for 3-digit). Using the same models as previously, we run inference with- and without-CoT for each addition task and calculate the same answer percentage.
@@ 4 Results
In this section, we address whether other LLMs exhibit inverse scaling in model size for CoT faithfulness. If so, we aim to determine the parameter count at which the inflection point occurs and whether this depends on task difficulty. Figure 2 shows how unfaithfulness (left), normalized unfaithfulness (center), and accuracy (right) change with different model sizes. We refer the reader to Tables 6-8 in the Appendix for a more detailed breakdown of these results.
Figure 3: Lanham et al. (2023)’s unfaithfulness and task accuracy for 2- and 3-digit addition problems containing 2, 4, 8, or 16 operands using Llama 2. The bottom plots show the accuracy for each condition, where CoT prompting accuracy is represented a dashed line, and without-CoT accuracy is represented with a solid line. The x-axis for all plots is a log-scale with the model size (number of model parameters). For both tasks the optimally faithful model according to the metric occurs at 13B; however, this might be due to the sparse nature of the x-axis
@@
Yes. In Figure 2, the Pythia DPO models exhibit a V-shaped pattern of CoT unfaithfulness similar to that identified by Lanham et al. (2023) in Figure 1 which flattens under the normalized metric. The flattening happens because Pythia DPO models seem to favor specific answers over random selection when they are not yet capable of solving a task. [^6] Appendix Figure 8 shows this tendency. Moreover, if we focus on the accuracy lines that are one standard deviation below the mean in Figure 2, we observe that only Llama 2 models and FLAN-UL2 ever become better than a random baseline for all benchmarks. [^7] A baseline that selects an answer choice at random ranges 20–25% depending on the benchmark. Accuracy results for individual benchmarks are found in Appendix Figure 10. We see inverse scaling in the Llama 2 unfaithfulness scores for both the unnormalized and normalized metrics; CoTs become less faithful as the models get bigger. These results support the takeaway of Lanham et al. (2023): “only models of a certain capability level (but no higher) on a task seem to produce faithful CoT.”
@@
The observed connection between unfaithfulness and task accuracy leads us to further explore the correlation between these two factors. In our accuracy–unfaithfulness plots, shown in Figure 4, we discover a high correlation in the normalized condition. We contend that this strikingly clear correlation between normalized unfaithfulness and accuracy with CoTs suggests a simplification of the intricate concept of measuring reasoning faithfulness that is too reductive.
@@
No. Lanham et al. (2023) find that inverse scaling only begins at 13B parameters. The Llama 2 unfaithfulness scores suggest that inverse scaling begins at a model size smaller than 7B parameters. The Pythia DPO unfaithfulness scores exhibit a V-shape similar to the one found by Lanham et al. (2023), however, it indicates inverse scaling begins at 1B parameters. After normalization, the Pythia DPO model exhibits a much shallower V-shape.
Figure 4: Task accuracy vs. Lanham et al. (2023)’s unfaithfulness metric on the left and task accuracy vs. normalized unfaithfulness on the right. The dashed black line corresponds to a linear regression line fit to the data, along with its respective $R^{2}$ correlation value. The correlation between accuracy and unfaithfulness is minimal ($R^{2}=0.025$), but strong ($R^{2}=0.740$) between accuracy and normalized unfaithfulness.
@@
No. To answer this, we consider Figure 3, which shows the Llama 2 models’ performance on the addition task. Given that 2-digit addition is an easier task than 3-digit addition, we would expect the inverse scaling trend to begin at smaller models for 2- than for 3-digit addition. We do not see this pattern in the Llama 2 unfaithfulness scores, where both settings point to the 13B model having the most faithful CoTs. The discrepancy in our findings compared to Lanham et al. (2023) may be an issue about granularity, where our three Llama 2 model sizes do not provide enough detail to see a difference between the addition conditions. Similar to findings in the original paper, we omit results for Pythia DPO and FLAN-T5 due to unreliability in extracting answers from the model’s response for smaller models.
@@ 5 Discussion and Conclusions
In §4, we demonstrate that high unfaithfulness scores by models performing no better than random do not necessarily stem from genuinely unfaithful CoTs but due to the models’ tendency to prefer specific answers. Thus, a more accurate understanding of CoTs faithfulness can be obtained by computing the normalized version of Lanham et al. (2023)’s unfaithfulness measurement that we use in this work. While it is possible that the smaller models used by Lanham et al. (2023) do not exhibit such a bias, which we cannot test directly since they are proprietary, its existence in the models we test highlights a potential flaw in the unnormalized metric.
@@
Moreover, after normalization, we find a strong linear correlation between task accuracy and unfaithfulness of generated CoTs. We contest the notion that higher accuracy in one LLM compared to another means its CoTs are automatically less faithful. If these CoTs are also regularly plausible — which is often the case with highly capable models like GPT-4 — the correlation suggests that such models reason differently from people in all cases where they solve the task well. Moreover, the notable change in the perceived unfaithfulness of the Pythia DPO 70M model, simply by accounting for sensitivity to answer choice order, raises concerns about the sensitivity of this measurement. Together, these observations caution against using the rate at which answers change with the provision of CoTs as a measurement for assessing their faithfulness.
6 Acknowledgments
We would like to thank the anonymous reviewers for their helpful feedback which led to eliminating a potential confounding variable in our experiments and a clearer presentation of our findings. We also thank the members of the UtahNLP group for valuable insights and discussion. The support and resources from the Center for High Performance Computing at the University of Utah are gratefully acknowledged.